OpenStudy (anonymous):

how do I add this program to the body of a for loop so that it reoccur monthly for one year?

OpenStudy (anonymous):

##OB = Outstanding Balance ##AIR = Annual Interest Rate ##MMPP = Minimum Monthly Payments Percentage ##MMP = Minimum Mnthly Payments ##IP = Interest Paid ##PP = Principal Paid ##RB = Remaining Balance OB=int(raw_input("Enter Outstanding Balance:")) AIR=float(raw_input("Enter Interest Rate(decimal):")) MMPP=float(raw_input("Enter Minimum Monthly Payment Percent(decimal):")) print "Month:", ##for each month of the year ##calculate minimum monthly payment ##??ad this amount to the total paid variable MMP = MMPP*OB ##print the min monthly payment print "Minimum Monthly Payment:",MMP ##calculate interest paid IP =(0.18/12.0)*OB ##Print Interest Paid print "Interest Paid:",IP ##Calculate Principal Paid, PP = Principal Paid PP = MMP-IP ##print Principal Paid print "Principal Paid:",PP ##Remaining Balanace RB = OB-PP ##print Remaining Balance print "Remaining Balance:",RB ##print total paid

OpenStudy (turingtest):

use either a while loop, or for loop: ``` month = 1 while month <= 12: ## do stuff print month month += 1 ``` or ``` for month in range (1,13): ## do stuff print month ``` notice that on each iteration, the variable month is updated, so printing out that variable will give the correct number for the month you are on

OpenStudy (anonymous):

thanks so much looks like that should be it!